# Closing the Gap Between SGP4 and High-Precision Propagation via Differentiable Programming

### 27 Feb 2024 v3

@@ 4.2 Experiments
This section presents experiments conducted to assess the accuracy of the ML-dSGP4 propagator and compare it with the uncorrected SGP4 baseline and a neural network corrector that only adjusts SGP4 outputs [7].
@@
The dataset comprises SpaceX Starlink Two-Line Elements (TLEs) and ephemeris data available on Space-Track888https://www.space-track.org/, date of access: April 2023.. The ephemeris data includes a set of 1,519 Starlink satellites, in the period from the 18th of February 2023, at 11:38:42 UTC, to the 21st of February 2023, at 11:38:42 UTC, with a time resolution of 60 seconds. This data is provided by SpaceX using their numerical orbital propagator. We also collect the TLE data for the same objects during the period of the 17th and 18th of February, before the date of the first available ephemeris. We then use dSGP4 to propagate those TLEs at the same times as the available ephemeris data, and we treat each TLE as a separate observation. This generates a dataset of as many entries as the number of TLEs for each object, times the number of time stamps in the ephemeris data, multiplied by the total number of satellites: that is, a dataset with 6,563,599 rows and 6 columns (three for the position vector and three for the velocity vector).
@@
We split the training, validation, and test sets at the TLE level, making sure that the data propagated from a certain TLE belongs to the same set and is not shared among multiple sets: this is to avoid the model essentially learning to interpolate the data. We use 69% of the dataset as training, 16% as validation, and 15% as testing. Moreover, we make sure that the test set covers ranges of altitude that are unseen in the training and validation set: this resulted in 225 Starlink satellites below 539 km (while all the other satellites above that altitude have been used in training and validation). We train the model with the architecture shown on the right side of Figure 2, by using two neural networks (each with three hidden layers of 35 neurons each), using 30 epochs, a learning rate of 3, mean squared error (MSE) as a loss, and Adam as optimizer. Moreover, we also learn the correction factors for all elements of the inputs and outputs of dSGP4 (this is a set of 12 parameters in total). The total number of learnable parameters was 3,454.
@@
We compare this model with a neural network output corrector that uses slightly more learnable parameters (i.e., 3596 parameters: to have a fair comparison): resulting from a feed-forward neural network with four hidden layers of size 32, plus the six correction factors for the output corrections. In both this and the ML-dSGP4 case, the networks only learn corrections to the given inputs and/or outputs.
@@
In Figure 3, we show the loss of the training, validation, and test data as a function of epochs for both the ML-dSGP4 and the neural network corrector model. As we observe, in all three datasets, both models manage to outperform the SGP4 baseline model, and ML-dSGP4 manages to achieve a higher accuracy than the neural network corrector, thereby confirming the added advantages of the differentiable setup that can also learn input corrections.
@@
In Table 2, we show the final losses of both models, compared to the baseline results: to produce the table, the best models at validation were chosen for the neural network corrector and ML-dSGP4.
@@
As already mentioned, the test set was in a set of altitudes (and therefore also dynamics) completely unseen during training or validation. Nonetheless, the ML-dSGP4 model manages to achieve accuracies in the test set that are in the same range of values as the validation ones: this confirms that the model successfully starts to learn a differential correction between the low-accuracy SGP4 and the high-precision orbital propagation model. This paves the way for a mixed strategy where low-precision propagation with dSGP4 can be combined with high-precision data (or orbital propagation data) to learn correction models that then enable to perform fast and more accurate predictions of orbital states. Besides, this is also combined with the possibility of performing these computations in batches and leveraging parallelization via CPU or GPU. Finally, in this experiment we did not learn any internal dSGP4 model correction factors: however, the same framework can also be leveraged to allow dSGP4 internal parameters to be learned, thereby providing further prediction accuracy enhancements.
@@
The objective of this experiment is to show that even with only three days of ephemeris data, the ML-dSGP4 model can be leveraged to increase the accuracy of the standard SGP4 propagator while retaining its computational advantages. Feeding more ephemeris data, the model can further enhance its accuracy, potentially starting to be competitive with high-precision numerical propagators in terms of precision, while retaining the advantages of general perturbation techniques in terms of speed. This might also be very appealing for operators of constellations that possess a large amount of high-precision data.
@@ 5.1 Methods
In the aforementioned works, either finite differences or analytical schemes are proposed as a means to find the Jacobian matrix needed for solving the least squares problem. However, in our case, we will show that this is not necessary, since the differentiability of dSGP4 enables the computation of those quantities via AD, which turns out to be very advantageous, as it avoids extra computation and/or additional errors.
@@
where are the observation weights. The modeled states at the observed times , can be found via the nonlinear propagation of at , until , using dSGP4:
It is straightforward to observe that the above problem is equivalent to a weighted least-squares problem, whose minimum can be sought starting from an initial guess and setting up an update rule that depends on the Jacobian of the modeled states at observation times, with respect to the state at the target time [4].
@@
Owing to the differentiability of dSGP4, the partial derivative terms in Eq. (7) can now be exactly computed via automatic differentiation. In case more than one observations are available, then the , matrices and vector can be found by stacking each observation: , . In case no weights are assigned for the observations, then the weight matrix corresponds to the identity matrix.
@@ 5.2 Experiments
Using the procedure highlighted in Section 5, we setup an experiment to show the usage of dSGP4 in an orbit determination problem. In our analysis, we assume to have a set of TLEs as observations, and we use SGP4 as a nonlinear orbital propagation model: our goal is to set up the problem as a nonlinear least squares problem, in which we want to find a TLE for a given target epoch that best fits the observations, using the SGP4 model. This experiment is also documented in one of the tutorials of the open-source package. As we already mentioned, three essential components of satellite orbit determination are a set of observations, a target time at which we desire to extract the satellites’ state, and an initial guess to setup the iterative procedure. We download the set of TLEs corresponding to Sentinel-2A satellite in the period from 1st of June to 8th of June 2022, from the Space-Track999https://www.space-track.org/, date of access: April 2023. public catalog. This corresponds to a set of 26 TLEs, with a maximum time difference of about 10 hours, and a minimum time difference of about 3 hours: we use these TLEs as observations. Then, we choose the target epoch, at which we would like to extract the best fit for the TLE, one day after the final observation, which corresponds to the 8th of June 2022, at 21:25:48.337248 UTC. The initial guess for the TLE elements is found from the average state of the 26 TLEs propagated at the target epoch via dSGP4. This corresponds to the following initial state: km, km/s, which results in the nominal target TLE (i.e., ) shown in Tab. 3. By propagating the nominal TLE at the observation times and comparing the norm of the residuals w.r.t. the actual observation, we get a total residual norm in the state of 2915.6622, with residual norms ranging from a maximum of 900 km in position and 1 km/s in velocity, to a minimum of 100 km in position and 0.1 km/s in velocity for the closest in time. We use this nominal TLE as an initial guess for the differential corrector. After convergence, the TLE in Tab. 4 is found, with a total residual norm in the state (norm of the sum of all residuals) of 1.2112, which corresponds to residuals that range from a minimum of 12.6 m in position and 0.5 m/s in velocity, and to a maximum of 513.6 m in position and 1.4 m/s in velocity. The difference between this example and previous satellite orbit determination applications shown in literature is that in this case, we used dSGP4 automatic differentiation to precisely compute the partial derivatives needed to build the matrix of Eq. (7), therefore avoiding the local truncation errors due to finite differencing [21].
@@ 6 Conclusions
In this work, we introduce dSGP4, a differentiable version of the SGP4 model implemented in PyTorch, representing a significant advancement in space situational awareness. Our research work yields several considerations.
@@
Moreover, we discussed the seamless integration of dSGP4 with state-of-the-art machine learning backpropagation algorithms. This opens new avenues for orbital propagation models. We therefore introduce ML-dSGP4, a machine learning SGP4 model to enhance SGP4 propagations to better match ephemeris and/or high-precision propagation data. During training, both dSGP4 parameters and neural network weights can be jointly optimized backpropagating through both the dSGP4 model and the neural networks, enhancing prediction accuracy to align more closely with observational data, such as ephemeris data. In this way, one can maintain the same SGP4 model efficiency and speed, while improving its accuracy and enabling parallelization over CPU and GPU. We show how the new propagation model is able to outperform both the SGP4 baseline and black-box neural network correctors on a sample of around 1,500 Starlink satellites.
Furthermore, the PyTorch implementation of dSGP4 facilitates straightforward parallelization on both CPU and GPUs, thereby allowing the processing of batches of TLE data at the same time. This acceleration holds promise for applications that display bottlenecks in terms of computational costs, such as many-to-many conjunction screening in an increasingly congested low-Earth orbit.
Throughout our paper, we present several experiments that delve into each of these applications, providing quantitative results that underscore the practical significance and versatility of dSGP4 in advancing space situational awareness and orbital prediction. Furthermore, we also release the dSGP4 code and associated tests and tutorials (which also go through the experiments in this paper) as an open-source Github repository under the ESA organization.

### 07 Mar 2024 v4

@@ 4.2 Experiments
This section presents experiments conducted to assess the accuracy of the ML-SGP4 propagator and compare it with the uncorrected SGP4 baseline and a neural network corrector that only adjusts SGP4 outputs [7].
@@
The dataset comprises SpaceX Starlink Two-Line Elements (TLEs) and ephemeris data available on Space-Track888https://www.space-track.org/, date of access: April 2023.. The ephemeris data includes a set of 1,519 Starlink satellites, in the period from the 18th of February 2023, at 11:38:42 UTC, to the 21st of February 2023, at 11:38:42 UTC, with a time resolution of 60 seconds. This data is provided by SpaceX using their numerical orbital propagator. We also collect the TLE data for the same objects during the period of the 17th and 18th of February, before the date of the first available ephemeris. We then use SGP4 to propagate those TLEs at the same times as the available ephemeris data, and we treat each TLE as a separate observation. This generates a dataset of as many entries as the number of TLEs for each object, times the number of time stamps in the ephemeris data, multiplied by the total number of satellites: that is, a dataset with about 6.5 million rows and 6 columns (three for the position vector and three for the velocity vector).
@@
We split the training, validation, and test sets at the TLE level, making sure that the data propagated from a certain TLE belongs to the same set and is not shared among multiple sets: this is to avoid the model essentially learning to interpolate the data. We use 69% of the dataset as training, 16% as validation, and 15% as testing. Moreover, we make sure that the test set covers ranges of altitude that are unseen in the training and validation set: this resulted in 225 Starlink satellites below 539 km (while all the other satellites above that altitude have been used in training and validation). We train the model with the architecture shown on the right side of Figure 2, by using two neural networks (each with three hidden layers of 35 neurons each), using 30 epochs, a learning rate of 3, mean squared error (MSE) as a loss, and Adam as optimizer. The total number of learnable parameters was 3,454.
@@
To have a fair comparison, we compare this model with a neural network output corrector that uses a similar number of learnable parameters (i.e., 3596): resulting from a feed-forward neural network with four hidden layers of size 32, plus the six correction factors for the output corrections. In both this and the ML-SGP4 case, the networks only learn corrections to the given inputs and/or outputs.
@@
In Figure 4, we show the loss of the training, validation, and test data as a function of epochs for both the ML-SGP4 and the neural network corrector model. As we observe, in all three datasets, both models manage to outperform the SGP4 baseline model, and ML-SGP4 manages to achieve a higher accuracy than the neural network corrector, thereby confirming the added advantages of the differentiable setup that can also learn input corrections.
@@
In Table 2, we show the final losses of both models, compared to the baseline results: to produce the table, the best models at validation were chosen for the neural network corrector and ML-SGP4.
@@
As already mentioned, the test set was in a set of altitudes (and therefore also dynamics) completely unseen during training or validation. Nonetheless, the ML-SGP4 model manages to achieve accuracies in the test set that are in the same range of values as the validation ones: this confirms that the model successfully starts to learn a differential correction between the low-accuracy SGP4 and the high-precision orbital propagation model. This paves the way for a mixed strategy where low-precision propagation with SGP4 can be combined with observed or simulated data to learn correction models that then enable to perform fast and more accurate predictions of orbital states. Besides, this is also combined with the possibility of performing these computations in batches and leveraging parallelization via CPU or GPU. Finally, in this experiment we did not learn any internal SGP4 model correction factors: however, the same framework can also be leveraged to allow SGP4 internal parameters to be learned, thereby providing further prediction accuracy enhancements.
@@
The objective of this experiment is to show that even with only three days of ephemeris data, the ML-SGP4 model can be leveraged to increase the accuracy of the standard SGP4 propagator while retaining its computational advantages. Feeding more ephemeris data, the model can further enhance its accuracy, potentially starting to be competitive with high-precision numerical propagators in terms of precision, while retaining the advantages of general perturbation techniques in terms of speed. This might also be very appealing for operators of constellations that possess a large amount of high-precision data.
@@ 5.1 Methods
In the aforementioned works, either finite differences or analytical schemes are proposed as a means to find the Jacobian matrix needed for solving the least squares problem. However, in our case, we will show that this is not necessary, since the differentiability of SGP4 enables the computation of those quantities via AD, which turns out to be very advantageous, as it avoids extra computation and/or additional errors.
@@
where are the observation weights. The modeled states at the observed times , can be found via the nonlinear propagation of at , until , using SGP4:
The above problem is equivalent to a weighted least-squares problem, whose minimum can be sought starting from an initial guess and setting up an update rule that depends on the Jacobian of the modeled states at observation times, with respect to the state at the target time [4].
@@
Owing to the differentiability of SGP4, the partial derivative terms in Eq. (7) can now be exactly computed via automatic differentiation. In case more than one observations are available, then the , matrices and vector can be found by stacking each observation: , . In case no weights are assigned for the observations, then the weight matrix corresponds to the identity matrix.
@@ 5.2 Experiments
Using the procedure highlighted in Section 5, we setup an experiment to show the usage of SGP4 in an orbit determination problem. In our analysis, we assume to have a set of TLEs as observations, and we use SGP4 as a nonlinear orbital propagation model: our goal is to set up the problem as a nonlinear least squares problem, in which we want to find a TLE for a given target epoch that best fits the observations, using the SGP4 model. This experiment is also documented in one of the tutorials of the open-source package. As we already mentioned, three essential components of satellite orbit determination are a set of observations, a target time at which we desire to extract the satellites’ state, and an initial guess to setup the iterative procedure. We download the set of TLEs corresponding to Sentinel-2A satellite in the period from 1st of June to 8th of June 2022, from the Space-Track999https://www.space-track.org/, date of access: April 2023. public catalog. This corresponds to a set of 26 TLEs, with a maximum time difference of about 10 hours, and a minimum time difference of about 3 hours: we use these TLEs as observations. Then, we choose the target epoch, at which we would like to extract the best fit for the TLE, one day after the final observation, which corresponds to the 8th of June 2022, at 21:25:48.337248 UTC. The initial guess for the TLE elements is found from the average state of the 26 TLEs propagated at the target epoch via SGP4. This corresponds to the following initial state: km, km/s, which results in the nominal target TLE (i.e., ) shown in Tab. 3. By propagating the nominal TLE at the observation times and comparing the norm of the residuals w.r.t. the actual observation, we get a total residual norm in the state of 2915.6622, with residual norms ranging from a maximum of 900 km in position and 1 km/s in velocity, to a minimum of 100 km in position and 0.1 km/s in velocity for the closest in time. We use this nominal TLE as an initial guess for the differential corrector. After convergence, the TLE in Tab. 4 is found, with a total residual norm in the state (norm of the sum of all residuals) of 1.2112, which corresponds to residuals that range from a minimum of 12.6 m in position and 0.5 m/s in velocity, and to a maximum of 513.6 m in position and 1.4 m/s in velocity. The difference between this example and previous satellite orbit determination applications shown in literature is that in this case, we used SGP4 automatic differentiation to precisely compute the partial derivatives needed to build the matrix of Eq. (7), therefore avoiding the local truncation errors due to finite differencing [21].
@@ 6 Conclusions
In this work, we introduce SGP4, a differentiable version of the SGP4 model implemented in PyTorch, representing a significant advancement in space situational awareness. Our research work yields several considerations.
@@
Moreover, we discussed the seamless integration of SGP4 with state-of-the-art machine learning backpropagation algorithms. This opens new avenues for orbital propagation models. We, therefore, introduce ML-SGP4, a machine learning SGP4 model to enhance SGP4 propagations to better match accurate simulated or observed data. During training, both SGP4 parameters and neural network parameters can be jointly optimized backpropagating through both the SGP4 model and the neural networks, enhancing prediction accuracy to align more closely with the data. In this way, one can maintain the same SGP4 model efficiency and speed, while improving its accuracy and enabling parallelization over CPU and GPU. We show how the new propagation model is able to outperform both the SGP4 baseline and black-box neural network correctors on a sample of around 1,500 Starlink satellites.
Furthermore, the PyTorch implementation of SGP4 facilitates straightforward parallelization on both CPU and GPUs, thereby allowing the processing of batches of TLE data at the same time. This acceleration holds promise for applications that display bottlenecks in terms of computational costs, such as many-to-many conjunction screening in an increasingly congested low-Earth orbit.
Throughout our paper, we present several experiments that delve into each of these applications, providing quantitative results that underscore the practical significance and versatility of SGP4 in advancing space situational awareness and orbital prediction. Furthermore, we also release the SGP4 code and associated tests and tutorials as an open-source Github repository under the ESA organization.
7 Acknowledgments
We gratefully acknowledge Dr. T.S. Kelso for sharing his expertise and providing valuable suggestions, particularly regarding the validation of our work.